# Predicting post-release defects with knowledge units (KUs) of programming languages: an empirical study

### v1

@@ 5.4.3 Findings
Table 5: P-value, effect size, and median AUC value of KUCLS, CC, and KUCLS+CC defect classification models.
Statistical Test
Median AUC KUCLS+CC vs KUCLS KUCLS+CC vs CC
Release KUCLS+CC KUCLS CC P-value Effect Size P-value Effect Size
activemq-5.0.0 0.91 0.82 0.89 <0.05 L(1.0) <0.05 L(0.82)
activemq-5.1.0 0.85 0.83 0.82 <0.05 M(0.46) <0.05 L(0.56)
activemq-5.2.0 0.87 0.83 0.85 <0.05 L(0.94) <0.05 L(0.69)
activemq-5.3.0 0.86 0.82 0.83 <0.05 L(0.85) <0.05 L(0.64)
activemq-5.8.0 0.88 0.85 0.86 <0.05 L(0.90) <0.05 L(0.80)
derby-10.2.1.6 0.88 0.74 0.85 <0.05 L(1.0) <0.05 L(0.88)
derby-10.3.1.4 0.88 0.84 0.85 <0.05 L(0.97) <0.05 L(0.90)
derby-10.5.1.1 0.87 0.82 0.84 <0.05 L(0.95) <0.05 L(0.77)
groovy-1.5.7 0.83 0.82 0.79 <0.05 S(0.17) <0.05 S(0.17)
groovy-1.6.BETA_1 0.90 0.82 0.87 <0.05 L(0.88) <0.05 L(0.54)
groovy-1.6.BETA_2 0.94 0.85 0.91 <0.05 L(0.90) <0.05 M(0.4)
hbase-0.94.0 0.86 0.72 0.84 <0.05 L(1.0) <0.05 M(0.45)
hbase-0.95.0 0.82 0.72 0.81 <0.05 L(1.0) <0.05 L(0.48)
hbase-0.95.2 0.73 0.71 0.71 <0.05 L(0.97) <0.05 L(0.54)
hive-0.10.0 0.86 0.72 0.83 <0.05 L(1.0) <0.05 L(0.75)
hive-0.12.0 0.86 0.80 0.84 <0.05 L(1.0) <0.05 L(0.54)
hive-0.9.0 0.91 0.79 0.9 <0.05 L(1.0) <0.05 L(0.53)
jruby-1.1 0.88 0.86 0.85 <0.05 S(0.25) <0.05 M(0.44)
jruby-1.4.0 0.86 0.83 0.84 <0.05 L(0.56) <0.05 L(0.49)
jruby-1.5.0 0.87 0.84 0.86 <0.05 L(0.59) <0.05 M(0.41)
jruby-1.7.0.preview1 0.86 0.84 0.83 <0.05 L(0.59) <0.05 L(0.74)
lucene-2.3.0 0.94 0.85 0.91 <0.05 L(1.0) <0.05 L(0.83)
lucene-2.9.0 0.84 0.79 0.82 <0.05 L(0.96) <0.05 L(0.66)
lucene-3.0.0 0.92 0.91 0.9 <0.05 S(0.32) <0.05 L(0.62)
lucene-3.1.0 0.79 0.74 0.76 <0.05 L(0.91) <0.05 M(0.44)
wicket-1.3.0-beta2 0.85 0.81 0.82 <0.05 L(0.76) <0.05 L(0.61)
wicket-1.3.0-incubating-beta-1 0.88 0.83 0.87 <0.05 L(0.82) <0.05 M(0.44)
wicket-1.5.3 0.82 0.81 0.79 <0.05 S(0.24) <0.05 L(0.65)
Median 0.87 0.82 0.84
@@
Observation 9) Combining KU-based features with traditional code metrics results in a higher performing model for predicting post-release defects. Figure 9 depicts the performance of the studied models. We observe that the combined model, KUCLS+CC, is the top performing model, surpassing all the studied models including the individual KUCLS and CC for predicting post-release defects. The median AUC of the KUCLS+CC is 0.87, representing a normalized improvement of 16.8% over the CC, 27.8% over the KUCLS and 35% over the CC_PROD. We further examine the performance of KUCLS+CC for individual releases. Table 5 presents the P-value, effect size, and median AUC for the KUCLS+CC, CC and KUCLS models for each studied release. We observe that the KUCLS+CC models outperform the CC and the KUCLS in all 28 studied releases, with the majority showing a large effect size (highlighted in green in the columns “KUCLS+CC vs KUCLS” and “KUCLS+CC vs CC”). We also calculate the normalized AUC improvement for each of the studied releases. The normalized AUC improvement of the KUCLS+CC over the CC ranges from 4.9% to 33.3%, while the improvement over the KUCLS ranges from 5.6% to 59.9%. We thus conclude that the the KUCLS+CC demonstrates superior performance for predicting post-release defects, making it more effective than either individual models.
@@
Figure 10: The rank distribution of features of the KUCLS+CC for defect classification. The number inside the square brackets indicates the final rank of the feature after applying Scott-Knott ESD for the second time. The KUs are highlighted with the blue colored font and traditional code metrics are highlighted with black colored font.
@@
Observation 10) DDEV is the top ranked important feature for predicting post-release defects while Inheritance and Method Encapsulation of the KU-based features share the rank as the third most important features for the KUCLS+CC. Figure 10 presents the rank distribution of each feature (final rank is shown in brackets) of the KUCLS+CC across the studied releases. The top two important features are the process metrics: DDEV (i.e., the number of distinct developers) is ranked first, followed by Added Lines (the normalized count of lines added to the module) as the second most important feature. We then have three features that are ranked third most important features. Two of these features (Inheritance and Method Encapsulation) are the KU-based features and the other one is the product metric (i.e., ConntClassCoupled). The inclusion of Inheritance and Method Encapsulation KUs among the top three features highlights the significant contribution of KU-based features to the performance of the KUCLS+CC in predicting post-release defects.

### v2

@@ 4.3 RQ3: Can KUM be made more accurate by combining its features with those from TM?
Table 4: The effect size and median AUC value of KUM, TM, and KUM+TM defect classification models.
Median AUC Statistical Test (Effect Size)
Release KUM+TM KUM TM KUM+TM vs KUM KUM+TM vs TM
activemq-5.0.0 0.92 0.83 0.89 L(1.00)* L(0.94)*
activemq-5.1.0 0.86 0.83 0.82 L(0.68)* L(0.81)*
activemq-5.2.0 0.89 0.83 0.85 L(0.98)* L(0.85)*
activemq-5.3.0 0.87 0.82 0.84 L(0.97)* L(0.81)*
activemq-5.8.0 0.89 0.85 0.86 L(0.97)* L(0.93)*
derby-10.2.1.6 0.89 0.74 0.86 L(1.00)* L(0.95)*
derby-10.3.1.4 0.90 0.84 0.86 L(1.00)* L(0.99)*
derby-10.5.1.1 0.89 0.83 0.85 L(1.00)* L(0.95)*
groovy-1.5.7 0.83 0.81 0.78 S(0.18)* M(0.40)*
groovy-1.6.BETA_1 0.90 0.82 0.87 L(0.81)* M(0.42)*
groovy-1.6.BETA_2 0.95 0.85 0.91 L(0.96)* L(0.74)*
hbase-0.94.0 0.86 0.71 0.83 L(1.00)* L(0.60)*
hbase-0.95.0 0.84 0.71 0.82 L(1.00)* L(0.68)*
hbase-0.95.2 0.76 0.67 0.73 L(1.00)* L(0.79)*
hive-0.10.0 0.88 0.72 0.85 L(1.00)* L(0.87)*
hive-0.12.0 0.88 0.72 0.85 L(1.00)* L(0.80)*
hive-0.9.0 0.94 0.79 0.91 L(1.00)* L(0.96)*
jruby-1.1 0.90 0.86 0.87 L(0.58)* L(0.63)*
jruby-1.4.0 0.88 0.83 0.85 L(0.84)* L(0.67)*
jruby-1.5.0 0.88 0.84 0.86 L(0.68)* L(0.54)*
jruby-1.7.0.preview1 0.89 0.84 0.85 L(0.93)* L(0.83)*
lucene-2.3.0 0.95 0.82 0.92 L(1.00)* L(0.91)*
lucene-2.9.0 0.85 0.78 0.82 L(0.99)* L(0.88)*
lucene-3.0.0 0.93 0.91 0.90 L(0.67)* L(0.91)*
lucene-3.1.0 0.80 0.73 0.75 L(0.96)* L(0.74)*
wicket-1.3.0-beta2 0.87 0.82 0.83 L(0.93)* L(0.81)*
wicket-1.3.0-incubating-beta-1 0.90 0.86 0.87 L(0.95)* L(0.80)*
wicket-1.5.3 0.83 0.81 0.79 M(0.41)* L(0.72)*
Median 0.89 0.82 0.85
@@
Findings. Observation 4) Combining KU features with traditional metrics results in a better performing model for predicting post-release defects. Figure 6 depicts the performance of the studied models. We observe that the combined model KUM+TM is the top performing model. The median AUC of KUM+TM is 0.89, representing a normalized improvement of 36.4% over TM and 63.6% over KUM. We further examine the performance of KUM+TM for individual releases (Table 4). We observe that KUM+TM models outperform TM and KUM in all 28 studied releases, with the majority showing a large effect size (highlighted in green in the columns “KUM+TM vs KUM” and “KUM+TM vs TM”). We also calculate the normalized AUC improvement for each of the studied releases. The normalized AUC improvement of KUM+TM over TM ranges from 11.1% to 44.4% (median 22.5%), while the improvement over KUM ranges from 10.5% to 72.2% (median 31.5%). We thus conclude that KUM+TM demonstrates superior performance for predicting post-release defects, making it more effective than either of the individual models.
@@
Figure 7: The rank distribution of features of KUM+TM. The number inside the square brackets indicates the final rank of the feature after applying Scott-Knott ESD for the second time. The KUs are highlighted with the blue colored font and traditional metrics are highlighted with black colored font.
@@
Observation 5) ADEV is the top ranked feature for predicting post-release defects. Method & Encapsulation is the highest ranked KU feature (fourth). Figure 7 presents the rank distribution of each feature (final rank is shown in brackets) of KUM+TM across the studied releases. The top two most important features are process metrics: ADEV (i.e., the number of active developers) and Added_Lines (the normalized count of lines added to the file). Major Commit, an ownership metric, is ranked third. We then have two features that are ranked as the forth most important feature. One of these features is Method & Encapsulation, which is a KU feature. The other KU features in the top ten set are Inheritance (sixth) and Exception Handling (ninth).